# Rasterizing Wireless Radiance Field via Deformable 2D Gaussian Splatting

### v2

@@ 4 Method
In this section, we present the SwiftWRF architecture for modeling WRF under single-sided transceiver mobility. We first formulate a 2D Gaussian representation to construct spatial spectra, then introduce a deformation mechanism to capture their dynamics during transceiver mobility. Finally, we describe a two-stage training protocol and key optimization strategies. The following subsections detail each component.
4.1 WRF Modeling via 2DGS
Under single-sided mobility, one transceiver is allowed to move, and a spatial spectrum is collected at each position where it stops. To model a collected spectrum $\mathbf{A}(\phi,\theta)$, we reformulate the 2D Gaussian formulation from Sec. 3.2 to the wireless domain and introduce a spatial rasterization process based on its original rasterization procedure.
2D Gaussian for WRF: In WRF, each 2D Gaussian primitive provides a localized representation of the signal, exhibiting a peak response at its center and decaying according to its associated Gaussian distribution. This distribution is parameterized by an angular center $\boldsymbol{\mu}_{n}=(\phi,\theta)$ and a covariance matrix $\boldsymbol{\Sigma}_{n}\in\mathbb{R}^{2\times 2}$. To model the WRF signal response, we introduce an attenuation factor $\delta_{n}\in\mathbb{R}$ to control amplitude decay, and a signal vector $\boldsymbol{\psi}_{n}\in\mathbb{R}^{2}$ to represent the real and imaginary parts of the complex-valued response. The angular center is normalized to lie within the valid spectrum range using a hyperbolic transformation:
@@ 4.1 WRF Modeling via 2DGS
Drawing inspiration from differentiable rasterization, we refer to this process as spectrum rasterization. The 2D Gaussian parameters are optimized iteratively through gradient backpropagation, guided by the loss between the predicted and ground-truth spectra. The entire pipeline is implemented in CUDA, with the spectrum domain partitioned into parallel tiles to fully exploit GPU acceleration.
4.2 Deforming 2DGS for WRF Modeling under Mobility
The spectrum rasterization described in Sec. 4.1 defines a one-to-one mapping between the spatial spectrum domain and the 2DGS domain. As the transceiver moves across positions $\boldsymbol{s}\in\mathcal{S}$, the resulting spectra $\{\mathbf{A}^{\boldsymbol{s}}\}_{\boldsymbol{s}\in\mathcal{S}}$ can be equivalently represented as a 2DGS video $\{\mathcal{G}^{\boldsymbol{s}}\}_{\boldsymbol{s}\in\mathcal{S}}$, where the frame $\mathcal{G}^{\boldsymbol{s}}$ consists of a set of 2D Gaussians
@@ 4.2 Deforming 2DGS for WRF Modeling under Mobility
However, modeling spectrum variations as residuals captures only discrete offsets at sampled positions $\boldsymbol{s}\in\mathcal{S}$, and fails to generalize to unseen locations $\boldsymbol{s}_{\text{new}}\notin\mathcal{S}$. To overcome this limitation, we introduce a continuous deformation field [^2] We use “deformation” to denote continuous variation and “residual” for discrete offsets., implemented as a lightweight MLP that maps any moving position to the corresponding primitive residuals (see Fig. 4 bottom left). Concretely, it predicts residuals for three components: First, $\Delta\boldsymbol{\mu}_{n}^{\boldsymbol{s}}$ represents the motion vector of each primitive, capturing translational shifts caused by transceiver mobility. Second, $\Delta\boldsymbol{\psi}_{n}^{\boldsymbol{s}}$ refines the signal response through a residual map that encodes fine-scale variations. Third, $\Delta\delta_{n}^{\boldsymbol{s}}$ adjusts the attenuation factor to account for phenomena such as path blockage or reappearance. The deformation MLP takes the positional encodings of each primitive center $\boldsymbol{\mu}$ and the moving position $\boldsymbol{s}$ as input, and predicts their residuals as
@@ 5 Synthetic spectra via ray-tracing
TABLE I: Hyperparameter setting.
@@
TABLE II: Comparison of methods on different scenes (Median PSNR↑, SSIM↑ and mean L1↓).
Method Conferenceroom (TX-moving) Bedroom (TX-moving) Office (TX-moving) NeRF2-s23
@@
Method Conferenceroom (RX-moving) Bedroom (RX-moving) Office (RX-moving) Average
@@ 6.3 Benchmark
Figure 8: Spectrum visualization on selected scenes (rendered at $90\times 360$); please zoom in to see details.
@@ 6.5 WRF reconstruction
1) Per-scene Evaluation: Table II summarizes the evaluation metrics comparing our method against all baselines across all evaluated scenes, which include the large-scale real-world dataset $\text{NeRF}^{2}-\text{s}23$ and the customized synthetic scenes. Overall, SwiftWRF achieves a PSNR improvement of 2.33–10.41 dB over competing methods, with consistent gains also observed in SSIM and l1 loss. On the large-scale $\text{NeRF}^{2}-\text{s}23$ dataset, SwiftWRF achieves a PSNR improvement of 2.99–4.03 dB over both $\text{NeRF}^{2}$ and WRF-GS, and especially outperforms the ray tracing method with a 63% increase in SSIM. This is attributed to the fact that ray tracing relies heavily on geometric priors, and the LiDAR-based scene reconstruction may be insufficiently accurate. Fig. 7 presents cumulative SSIM curves across all evaluated positions, demonstrating that our method outperforms all baselines by a substantial margin. On synthetic scenes with simpler layouts, SwiftWRF consistently achieves superior results, with a PSNR improvement ranging from 2.65 dB to 10.41 dB over all baselines. Notably, the comparable performance observed under both mobility scenarios suggests that SwiftWRF captures two distinct types of field variation: in the RX-moving case, the MLP likely models spatial variations across different receiving positions, whereas in the TX-moving case, it captures temporal variations at a fixed location.
Figure 9: Comparison of inference time for a single spectrum.
Figure 10: Evaluation on different training scales.
@@
TABLE III: Ablation study of deformation components.
$\checkmark$ $\checkmark$ $\checkmark$ 0% / 0%
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -19.10% / +86.44%
$\checkmark$ $\checkmark$ -15.73% / +69.49%
$\checkmark$ $\checkmark$ -1.22% / +1.69%
$\checkmark$ $\checkmark$ -3.37% / +13.56%
TABLE IV: Ablation study of training items.
[…]
N/A 0% / 0% 0%
w/o Stop Gradient +0.09% / +19.2% +53.9%
w/o Coarse Training +0.05% / -12.21% -0.06%
w/o Anneal Smoothing -0.17% / +2.72% -3.79%
w/o SSIM Loss -8.53% / +26.91% -6.11%
w/ L2 Loss -7.87% / +24.12% +12.13%
[…]
@@
6.6 Ablation Study
In this section, we conduct a comprehensive ablation study of the SwiftWRF design. To validate the effectiveness of each design choice, the ablation is organized into two parts: (1) deformation components and (2) training protocols. For (1), as shown in Table III, applying deformation to all components does not yield the best performance as initially expected. In particular, adding covariance deformation leads to an 86.44% increase in l1 loss, primarily due to the numerical instability it introduces by violating the lower triangular structure required for Cholesky decomposition. Excluding the covariance branch, all other deformation components contribute positively to performance, as verified through our step-wise ablation. Notably, position deformation proves to be the most critical, with its removal resulting in a 13.56% increase in l1 loss. Signal deformation has a relatively minor impact on reconstruction accuracy but still provides measurable gains in overall performance. These findings highlight the complementary roles of each component and confirm the optimality of our deformation design. For (2), the results reported in Table IV show that applying stop-gradient significantly reduces training time, with only a marginal drop in reconstruction performance. All other training configurations have relatively minor impact on training speed, except for the use of l2 loss, which introduces a 12.13% overhead. The results also highlight that SSIM regularization and anneal smoothing reduce l1 loss to 2.72% and 26.91%, respectively, demonstrating their effectiveness in mitigating overfitting.
@@ 7.1 AoA Prediction
Results: Fig. 12 presents the CDF of azimuth and pitch angle prediction error under four turbo learning mixture ratios: 10%, 30%, 50%, and 70% [^3] The mixture ratio refers to the proportion of ground-truth spectra in the training set. For example, a 10% ratio indicates that only 10% of the samples are ground truth, while the remaining 90% are generated spectra.. The purple curve labeled as GT (100%) represents the upper bound of prediction fidelity, obtained through the naïve training strategy described above. In contrast, the GT curves corresponding to each mixture ratio serve as theoretical lower bounds, as they retain only the specified proportion of the ground-truth training set. Fig. 12 uses red arrows to highlight the performance gains from SwiftWRF-generated spectra, with improvements of 18%, 21%, 16%, and 17% across the four ratios. As reported in Table V, we further assess SwiftWRF’s effectiveness by comparing its turbo learning performance against NeRF2 and WRF-GS under the same mixture ratios. Across all mixture ratios, SwiftWRF achieves the lowest median prediction errors in both pitch and azimuth. In particular, it outperforms NeRF2 and WRF-GS in azimuth prediction with relative improvements ranging from 8.7%–30.6% and 19.2%–60.7%, respectively.
7.2 RSSI Prediction
TABLE V: Median pitch and azimuth prediction errors (in degrees).
@@ 7.2 RSSI Prediction
Pitch ↓ Azimuth ↓ Pitch ↓ Azimuth ↓ Pitch ↓ Azimuth ↓
@@ 8 Conclusion
Limitation: While SwiftWRF enables efficient spectrum reconstruction, its WRF representations are currently trained offline. Incorporating online updates based on transceiver mobility could enable real-time channel estimation and enhance system adaptability. Additionally, although SwiftWRF delivers high-fidelity results in most cases, occasional synthesis failures indicate that further improvements in robustness and reconstruction quality remain necessary.

### v3

@@ 4 Method
In this section, we present the SwiftWRF architecture for modeling WRF under single-sided transceiver mobility. We first formulate a 2D Gaussian representation to construct the spatial spectrum, then introduce a deformation mechanism to capture their dynamics during transceiver mobility. Finally, we describe a two-stage training protocol and key optimization strategies. The following subsections detail each component.
4.1 Modeling WRF Spectrum via 2DGS
Under single-sided mobility, one transceiver is allowed to move, and a spatial spectrum of the WRF is collected at each position where it stops. To model a collected spectrum $\mathbf{A}(\phi,\theta)$, we reformulate the 2D Gaussian formulation from Sec. 3.2 to the wireless domain and introduce a spatial rasterization process based on its original rasterization procedure.
Wireless 2D Gaussian: In the wireless domain, each 2D Gaussian primitive provides a localized representation of the spectrum signal, exhibiting a peak response at its center and decaying according to its associated Gaussian distribution. This distribution is parameterized by an angular center $\boldsymbol{\mu}_{n}=(\phi,\theta)$ and a covariance matrix $\boldsymbol{\Sigma}_{n}\in\mathbb{R}^{2\times 2}$. To model the signal response, we introduce an attenuation factor $\delta_{n}\in\mathbb{R}$ to control amplitude decay, and a signal vector $\boldsymbol{\psi}_{n}\in\mathbb{R}^{2}$ to represent the real and imaginary parts of the complex-valued response. The angular center is normalized to lie within the valid spectrum range using a hyperbolic transformation
@@ 4.1 Modeling WRF Spectrum via 2DGS
Drawing inspiration from differentiable rasterization, we refer to this process as spectrum rasterization, as shown in Fig. 4. The 2D Gaussian parameters are optimized iteratively through gradient backpropagation, guided by the loss between the predicted and ground-truth spectra. The entire pipeline is implemented in CUDA, with the spectrum domain partitioned into parallel tiles to fully exploit GPU acceleration.
4.2 Deforming 2DGS for Mobility Modeling
Figure 5: 2DGS deformation for spectrum rendering under single-sided transceiver mobility.
The spectrum rasterization described in Sec. 4.1 defines a one-to-one mapping between the spatial spectrum domain and the 2DGS domain. As the transceiver moves across positions $\boldsymbol{s}\in\mathcal{S}$, the resulting spectra $\{\mathbf{A}^{\boldsymbol{s}}\}_{\boldsymbol{s}\in\mathcal{S}}$ can be equivalently represented as a 2DGS video $\{\mathcal{G}^{\boldsymbol{s}}\}_{\boldsymbol{s}\in\mathcal{S}}$, where each frame $\mathcal{G}^{\boldsymbol{s}}$ consists of a set of 2D Gaussians
@@ 4.2 Deforming 2DGS for Mobility Modeling
However, modeling spectrum variations as residuals captures only discrete offsets at sampled positions $\boldsymbol{s}\in\mathcal{S}$, and fails to generalize to unseen locations $\boldsymbol{s}_{\text{new}}\notin\mathcal{S}$. To overcome this limitation, we introduce a continuous deformation field [^1] We use “deformation” to denote continuous variation and “residual” for discrete offsets., implemented as a lightweight MLP that maps any moving position to the corresponding primitive residuals. Concretely, it predicts residuals for three components: First, $\Delta\boldsymbol{\mu}_{n}^{\boldsymbol{s}}$ represents the motion vector of each primitive, capturing translational shifts caused by transceiver mobility. Second, $\Delta\boldsymbol{\psi}_{n}^{\boldsymbol{s}}$ refines the signal response through a residual map that encodes fine-scale variations. Third, $\Delta\delta_{n}^{\boldsymbol{s}}$ adjusts the attenuation factor to account for phenomena such as path blockage or reappearance. The deformation MLP takes the positional encodings of each primitive center $\boldsymbol{\mu}$ and the moving position $\boldsymbol{s}$ as input, and predicts their residuals as
@@ 5 Synthetic spectra via ray-tracing
TABLE I: Hyperparameter setting
@@
TABLE II: Comparison of methods across different scenes, reported with median PSNR (↑), SSIM (↑), and mean L1 (↓). The best result is highlighted in dark gray, and the second best in light gray
Method Conferenceroom (moving TX) Bedroom (moving TX) Office (moving TX) NeRF2-s23
@@
Method Conferenceroom (moving RX) Bedroom (moving RX) Office (moving RX) Average
@@ 6.2 Dataset
Figure 8: Spectrum visualizations on selected scenes, including two positions from the NeRF²-23 public dataset and one position each from bedroom (RX-side moving), conference room (RX-side moving), and conference room (TX-side moving) in our dataset. Rendered at $90\times 360$); please zoom in to see details.
@@ 6.5 WRF reconstruction
TABLE III: Comparisons with recent GS-based methods on the NeRF2-s23 dataset
[…]
RFSPM N/A 4.2 N/A
GSpaRC 0.33 0.39 0.82
Ours 0.83 0.01 0.87
TABLE IV: Comparison of model storage and training time to convergence
Method Storage (MB) $\downarrow$ Training time (to convergence) $\downarrow$
NeRF2 8.00 5.8 h
WRF-GS 18.15 1.5 h
Ours 3.47 50 min
1) Per-scene Evaluation: Table II summarizes the evaluation metrics comparing our method against all baselines across all evaluated scenes, which include the large-scale real-world dataset $\text{NeRF}^{2}-\text{s}23$ and the customized synthetic scenes. Overall, SwiftWRF achieves a PSNR improvement of 2.33–10.41 dB over competing methods, with consistent gains also observed in SSIM and l1 loss. On the large-scale $\text{NeRF}^{2}-\text{s}23$ dataset, SwiftWRF achieves a PSNR improvement of 2.99–4.03 dB over both $\text{NeRF}^{2}$ and WRF-GS, and especially outperforms the ray tracing method with a 63% increase in SSIM. This is attributed to the fact that ray tracing relies heavily on geometric priors, and the LiDAR-based scene reconstruction may be insufficiently accurate. Fig. 9 presents cumulative SSIM curves across all evaluated positions, demonstrating that our method outperforms all baselines by a substantial margin. On synthetic scenes with simpler layouts, SwiftWRF consistently achieves superior results, with a PSNR improvement ranging from 2.65 dB to 10.41 dB over all baselines. Notably, the comparable performance observed under both mobility scenarios suggests that SwiftWRF captures two distinct types of field variation: in the RX-moving case, the MLP likely models spatial variations across different receiving positions, whereas in the TX-moving case, it captures temporal variations at a fixed location.We also compare our method with recent GS-based approaches [37, 25] on the public NeRF2-s23 dataset, as summarized in Table III. Overall, our method achieves substantially faster rendering while maintaining comparable or better reconstruction quality. Notably, we reach a median SSIM of 0.85 within 0.33 hours, already surpassing GSpaRC’s reported converged performance. Since RFSPM does not release its code, we report comparisons only against the metrics provided in their paper.
@@
Figure 12: Comparison of inference time for a single spectrum.
Figure 13: Evaluation on different training scales.
Figure 14: Wireless radiance field distribution visualization on Office and ConferenceRoom. Circles indicate a subset of training samples.
Figure 15: Toy example of 2DGS deformation. Left: a sample TX trajectory. Middle: per-Gaussian center deformation versus trajectory length. Right: deformation magnitude versus trajectory length.
3) Efficiency: We evaluate both the training and inference efficiency of a single-spectrum reconstruction across all methods on the $\text{NeRF}^{2}-\text{s23}$ dataset, with results summarized in Fig. 12 and Fig. 11 (a). All methods are executed on a single NVIDIA RTX 3090 GPU, except for ray tracing. As shown in Fig. 12, our method achieves a substantial performance advantage, reaching approximately 100k FPS—over 500× faster than WRF-GS and more than 20,000× faster than $\text{NeRF}^{2}$. In addition, our model consistently outperforms both WRF-GS and $\text{NeRF}^{2}$ during training, achieving around 20% higher median SSIM across varying rendering times (see Fig. 11 (a)). This efficiency arises from our compact 2D Gaussian formulation, which enables highly parallelized CUDA-based rasterization without incurring the overhead of volumetric rendering in $\text{NeRF}^{2}$ or the Jacobian-based projection computations required by WRF-GS. Beyond its fast speed, SwiftWRF also delivers superior compactness, requiring only 1202 MB of GPU memory compared with 1804 MB for WRF-GS and 8978 MB for $\text{NeRF}^{2}$.
@@
6) Storage: According to Table IV, our method is the most compact among NeRF2 and WRF-GS, reducing the model storage from 8.00 MB (NeRF2) and 18.15 MB (WRF-GS) to 3.47 MB, which corresponds to 56.6% and 80.9% less storage, respectively.
7) Resolution: We set the spectrum resolution to $90\times 360$ following NeRF2’s evaluation protocol, but our method is not limited to this setting. Thanks to the explicit primitive-based formulation of Gaussian splatting, it can scale to higher resolutions by increasing the number of Gaussians. We additionally evaluate on Bedroom at $0.5^{\circ}$ resolution ($180\times 720$), and observe effective reconstruction at this higher resolution. On the test set, we report an L1 loss of $0.039$ and a median SSIM of $0.898$. Training is extended slightly from 50 minutes to 70 minutes.
TABLE V: Ablation study of deformation components
$\checkmark$ $\checkmark$ $\checkmark$ 0.9040 / 0.0417
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.7313 / 0.0777
$\checkmark$ $\checkmark$ 0.7618 / 0.0707
$\checkmark$ $\checkmark$ 0.8930 / 0.0424
$\checkmark$ $\checkmark$ 0.8735 / 0.0474
TABLE VI: Ablation study of training items
[…]
N/A 0.9040 / 0.0417 50.0
w/o Stop Gradient 0.9048 / 0.0497 77.0
w/o Coarse Training 0.9045 / 0.0366 50.0
w/o Anneal Smoothing 0.9025 / 0.0428 48.0
w/o SSIM Loss 0.8269 / 0.0529 47.0
w/ L2 Loss 0.8329 / 0.0518 56.0
6.6 Visualization and analysis
To better illustrate the learned field and deformation behavior, we include two additional visualizations in the evaluation section. First, Fig. 14 reports the full-field distribution for Bedroom and ConferenceRoom. At each sampled position, we compute the RSSI by averaging the spectrum magnitude over both azimuth and elevation directions, and visualize the resulting ground-truth field together with the field reconstructed by our method. Second, Fig. 15 visualizes the deformation of 2D Gaussian primitives along a continuous TX trajectory. We extract a sample trajectory from the two scenes using a nearest-walk strategy and plot the per-Gaussian center deformation as well as the corresponding deformation magnitude across 60 trajectory steps. Notably, the deformation patterns vary locally along the trajectory, indicating that our position-conditioned deformation field models the residuals at different locations in a largely independent manner.
6.7 Ablation Study
In this section, we conduct a comprehensive ablation study of the SwiftWRF design. To validate the effectiveness of each design choice, the ablation is organized into two parts: (1) deformation components and (2) training protocols. For (1), as shown in Table V, applying deformation to all components does not yield the best performance as initially expected. In particular, adding covariance deformation leads to an 86.44% increase in l1 loss, primarily due to the numerical instability it introduces by violating the lower triangular structure required for Cholesky decomposition. Excluding the covariance branch, all other deformation components contribute positively to performance, as verified through our step-wise ablation. Notably, position deformation proves to be the most critical, with its removal resulting in a 13.56% increase in l1 loss. Signal deformation has a relatively minor impact on reconstruction accuracy but still provides measurable gains in overall performance. These findings highlight the complementary roles of each component and confirm the optimality of our deformation design. For (2), the results reported in Table VI show that applying stop-gradient significantly reduces training time, with only a marginal drop in reconstruction performance. All other training configurations have relatively minor impact on training speed, except for the use of l2 loss, which introduces a 12.13% overhead. The results also highlight that SSIM regularization and anneal smoothing reduce l1 loss to 2.72% and 26.91%, respectively, demonstrating their effectiveness in mitigating overfitting.
@@ 7.1 AoA Prediction
Results: Fig. 16 presents the CDF of azimuth and pitch angle prediction error under four turbo learning mixture ratios: 10%, 30%, 50%, and 70% [^2] The mixture ratio refers to the proportion of ground-truth spectra in the training set. For example, a 10% ratio indicates that only 10% of the samples are ground truth, while the remaining 90% are generated spectra.. The purple curve labeled as GT (100%) represents the upper bound of prediction fidelity, obtained through the naïve training strategy described above. In contrast, the GT curves corresponding to each mixture ratio serve as theoretical lower bounds, as they retain only the specified proportion of the ground-truth training set. Fig. 16 uses red arrows to highlight the performance gains from SwiftWRF-generated spectra, with improvements of 18%, 21%, 16%, and 17% across the four ratios. As reported in Table VII, we further assess SwiftWRF’s effectiveness by comparing its turbo learning performance against NeRF2 and WRF-GS under the same mixture ratios. Across all mixture ratios, SwiftWRF achieves the lowest median prediction errors in both pitch and azimuth. In particular, it outperforms NeRF2 and WRF-GS in azimuth prediction with relative improvements ranging from 8.7%–30.6% and 19.2%–60.7%, respectively.
TABLE VII: Median pitch and azimuth prediction errors (in degrees)
@@
pitch angle ↓ azimuth angle ↓ pitch angle ↓ azimuth angle ↓ pitch angle ↓ azimuth angle ↓
@@
7.2 RSSI Prediction
TABLE VIII: Median L1 loss (dB) comparison across methods
Method NeRF2 MRI SwiftWRF
Median L1 loss (dB)↓ 3.05 6.74 2.91
@@ 8 Conclusion
Discussion and Limitation: While SwiftWRF enables efficient spectrum reconstruction, our current model is trained offline and assumes a fixed environment during deployment. Incorporating lightweight online updates under transceiver mobility is a promising direction to further improve real-time channel estimation and system adaptability. Moreover, our current formulation models the measured spectrum outcomes rather than explicitly physical propagation effects such as visibility/path length or time-of-flight, and we keep the covariance updates fixed for stability. These design choices prioritize efficiency and stable optimization, but they may reduce physical interpretability. Although SwiftWRF achieves high fidelity in most evaluations, occasional synthesis failures and challenging transition regions suggest that further improvements in robustness and reconstruction quality are still needed. In addition, SwiftWRF does not currently support the NeRF2 channel-estimation setting, which requires predicting high-dimensional CSI matrices from paired uplink/downlink measurements; developing a more stable multi-channel parameterization and training strategy is left for future work.